# Discovering Artificial Viscosity Models for Discontinuous Galerkin approximation of Conservation Laws using Physics-Informed Machine Learning

### v1

@@ 5.3.3 Test case 9: Euler equation, Riemann problem
NN EV NN EV NN EV
$\epsilon$ 1.8409e+04 3.1180e+04 1.2387e+04 2.1667e+04 1.1578e+04 1.7845e+04
$\nabla\epsilon$ 2.0193e+04 2.9233e+04 7.0429e+04 9.1563e+04 1.3498e+05 1.6098e+05
$[\![\epsilon]\!]$ 5.5618e+03 4.5949e+03 2.3759e+03 1.7313e+03 1.4495e+03 9.8379e+02
o/u 7.4932e+01 1.7049e+02 6.5138e+01 1.3598e+02 3.9249e+01 7.5978e+01
@@
Since the problem is completed with periodic boundary conditions, we enlarge the domain to $\Omega=(-\frac{3}{2},\frac{3}{2})^{2}$ to avoid contaminations of the solution. The discretization is done with a structured grid with $7200,1800,450$ triangular elements and $k=1,3,5$, $\text{CFL}=0.05,0.18,0.21$, respectively. The results are reported in Figure 2324. A qualitative way to evaluate the model is to check the error done in representing the fine structures in $\boldsymbol{x}=[0,0]^{\top}$ and $\boldsymbol{x}=[0.4,0.4]^{\top}$. It is possible to notice that the NN model captures better the contact wave, especially for $k=3$. This is confirmed by a quantitative analysis of the errors, as shown in Table 8.
@@ 6 Conclusions
This algorithm also opens the possibility of training models by incorporating data that comes from the physical measurement of quantities of interest, such as the velocity or the pressure of a flow field. In future works, we aim to test that this integration effectively enhances model reliability by incorporating fundamental principles and constraints, leading to more accurate predictions and improved generalization to real-world scenarios. Moreover, we plan to apply the proposed algorithm to more domain-specific test cases such as the variational multiscale stabilization methods for the approximation of the incompressible Navier-Stokes equations.

### v2

@@ 5.3.3 Test case 9: Euler equation, Riemann problem
NN Ours EV NN [26] NN Ours EV NN [26] NN Ours EV
$\epsilon$ 1.8409e4 3.1180e4 1.4657e4 1.2387e4 2.1667e4 1.1576e4 1.1578e4 1.7845e4
$\nabla\epsilon$ 2.0193e4 2.9233e4 1.6856e4 7.0429e4 9.1563e4 6.6923e4 1.3498e5 1.6098e5
$[\![\epsilon]\!]$ 5.5618e3 4.5949e3 7.2912e3 2.3759e3 1.7313e3 4.4055e3 1.4495e3 9.8379e2
o/u 7.4932e1 1.7049e2 6.5821e2 6.5138e1 1.3598e2 5.0583e2 3.9249e1 7.5978e1
mv 5.911e-14 5.570e-14 5.656e-14 6.366e-14 6.480e-14 4.768e-14 7.731e-14 7.162e-14
@@
Since the problem is completed with periodic boundary conditions, we enlarge the domain to $\Omega=(-\frac{3}{2},\frac{3}{2})^{2}$ to avoid contaminations of the solution. The discretization is done with a structured grid with $7200,1800,450$ triangular elements and $k=1,3,5$, $\text{CFL}=0.05,0.18,0.21$, respectively. The Courant number is $C=0.1098,0.0443,0.0179$. We compare our NN with a tuned EV model and the NN of [16]. The results are reported in Figure 23. A qualitative way to evaluate the model is to check the error done in representing the fine structures in $\boldsymbol{x}=[0,0]^{\top}$ and $\boldsymbol{x}=[0.4,0.4]^{\top}$. It is possible to notice that the NN model captures better the contact wave, especially for $k=3$. This is confirmed by a quantitative analysis of the errors, as shown in Table 8. In this case, the model of [16] slightly outperforms our NN, producing more accurate results under several metrics. Our NN model injects viscosity greater than $10^{-8}$ in 3730, 1023, and 316 cells for $k=1,3,5,$ respectively.
@@ 6 Conclusions
This algorithm also opens the possibility of training models by incorporating data that comes from the physical measurement of quantities of interest, such as the velocity or the pressure of a flow field. In future works, we aim to test that this integration effectively enhances model reliability by incorporating fundamental principles and constraints, leading to more accurate predictions and improved generalization to real-world scenarios. For instance, integrating in the loss a regularization term based on the entropy production could potentially improve the model’s adherence to physical laws. Moreover, we plan to apply the proposed algorithm to more domain-specific test cases such as the variational multiscale stabilization methods for the approximation of the incompressible Navier-Stokes equations.